OpenStudy (anonymous):

How do I write these into standard quadratic form? y= 3(x-2)^2-5 f(x)= -2(x+4)^2+3

OpenStudy (mathmale):

it'd be good to look up "standard quadratic form." I'd suggest you do that. I'll drop you a hint: See http://www.mathsisfun.com/algebra/quadratic-equation.html In the first case, \[y=3(x-2)^2 - 5\] should be re-written by doing the following: 1. expand (x-2)^2. 2. Multiply each resulting term by 3. 3. Subtract 5 from the result. 4. combine like terms and re-arrange them in descending order by power of x.

OpenStudy (anonymous):

\[3x ^{2}-17\] ?!

OpenStudy (mathmale):

Would you mind showing how you obtain your answer, please? I'd be especially interested in seeing how you expand (x-2)^2.

OpenStudy (anonymous):

\[y=3(x-2)^{2}-5\] \[(x-2)^{2}=x ^{2}-4\] \[3(x ^{2}-4)=3x ^{2}-12\] \[3x ^{2}-12-5=3x ^{2}-17\]

OpenStudy (anonymous):

(x-2)^2 = x^2 -4x + 4

OpenStudy (anonymous):

How do you get that? @douglaswinslowcooper

OpenStudy (mathmale):

\[(x-2)^{2}=x ^{2}-4\]is incorrect. Please recall that (x-2)^2 signifies (x-2)(x-2). Try again, please.

OpenStudy (mathmale):

Important: Please show your work here.

OpenStudy (anonymous):

Do I do distributive property?

OpenStudy (mathmale):

That's certainly part of what you need to do. Many students refer to the "FOIL" method of multiplying binomials together.

OpenStudy (anonymous):

What's FOIL? I'm used to PEMDAS. (Parentheses, Exponents, Multiplying, Division, Addition, Subtraction)

OpenStudy (anonymous):

(x-2)(x-2) = x(x-2) - 2(x-2) = x^2-2x-2x+4=x^2-4x+4

OpenStudy (anonymous):

I don't get that.

OpenStudy (mathmale):

FOIL refers to First, Outer, Inner, Last. In (x-2)(x-2), the First terms are x and x. Multiplying them together, we get x^2. The Outer terms are x and -2. The Inner terms are -2 and x. The Last terms are -2 and -2. See this before?

OpenStudy (anonymous):

No.

OpenStudy (mathmale):

douglaswinslowcooper has a different method to multiply (x-2)(x-2), but his result is correct. In summary, you very much do need to know how to multiply binomials. Ask for further clarification if you need it.

OpenStudy (mathmale):

Do you have an algebra textbook?

OpenStudy (anonymous):

Yes.

OpenStudy (mathmale):

Look up "special products and factoring." Or look at http://www.wtamu.edu/academic/anns/mps/math/mathlab/int_algebra/int_alg_tut29_specfact.htm.